# Zero-shot Safety Prediction for Autonomous Robots with Foundation World Models

### v1

@@ III Foundation World Models
TABLE III: Safety prediction performance for the cart pole: F1 score and false positive rate. Gray rows indicate the use of additional data.
IV Experiments
IV-A Experimental Setup
We use simulation environments from OpenAI gym [10] as two cases in this paper: (i) a cart pole with a four-dimensional physical state and (ii) a lunar lander with an eight-dimensional physical state. For the cart pole, the safety threshold $\theta_{thre}$ is defined as the angle between the pole and the vertical line $\theta$ being less than $\pi/4$: $|\theta|<\pi/4$. For the lunar lander, safety means keeping the horizontal position $d_{x}$ of the lander within the landing range: $128<d_{x}<$ 12. For the cart pole, to evaluate safety from our predicted states, we calculate the angle instead of showing the pole’s centroid error. Given that the pole is upright most of the time, to get a more detailed evaluation of our models, we split the test experiment into two parts: (i) upright and (ii) falling.
As supervised learning baselines, we train two extra state prediction models: a Multilayer Perceptron (MLP) and a Long Short-Term Memory (LSTM). They are combined with our SAM-based representations. We also compare our approach to the original world model based on VAE representations and an MDN-LSTM state predictor. We test for 10,000 sequences with different input lengths $m$ and prediction horizons $k$ in each case study. In the results tables, we marked the results of the supervised models in grey because these models take 30,000 more sequences to train and therefore impose a significant data burden compared to our zero-shot foundation world models. This fact relaxes the expectations of the performance of our proposed models.
Method Input length F1 score FPR
k=10 k=20 k=30 k=40 k=50 k=60 k=10 k=20 k=30 k=40 k=50 k=60
SAM & MLP .$0$$0$$0$$0$$0$$0$3914 .2845 .2074 .1002 .0535 .0000
SAM & LSTM $m$=1 .$0$$0$$0$$0$$0$$0$0676 .0784 .0579 .0579 .0289 .0000
SAM & MLP .$0$$0$$0$$0$$0$$0$4456 .2692 .2161 .1320 .0400 .0000
SAM & LSTM $m$=2 .$0$$0$$0$$0$$0$$0$0685 .0802 .0536 .0606 .$0$$0$
SAM & Gemma .8850 .9038 .3000 1.000 .8651 .8329 .1441 .$0$$0$$0$0221 .0375
SAM & MLP .$0$$0$$0$$0$$0$$0$5473 .4475 .3860 .3093 .2871 .2889
SAM & LSTM $m$=4 .$0$$0$$0$$0$$0$$0$0825 .0837 .0775 .0597 .$0$$0$
SAM & MLP .$0$$0$$0$$0$$0$$0$$0$$0$$0$$0$$0$$0$
SAM & LSTM $m$=8 .$0$$0$$0$$0$$0$$0$0527 .0667 .0742 .0579 .0460 .0638
SAM & GPT 3.5 .9168 .9710 .6575 .8233 .3260 .6042 .0064 .0068 .$0$$0$$0$$0$
TABLE IV: Safety prediction performance for the lunar lander: F1 score and false positive rate. Gray rows indicate the use of additional data.
@@ IV-B Experimental Results
As depicted in Table I, the errors of all methods are relatively low on short prediction horizons. As the value of $k$ increases, the error generally increases, indicating that predicting states further into the future is more difficult. Specifically, it can be noted that GPT 3.5 and Gemma alternately exhibit relatively low errors in most settings, especially when predicting the falling status. Table II shows the error for the lunar lander, where the foundation models do not perform as well as supervised learning methods on short inputs. Our investigation suggests that this shortcoming is due to insufficiently long input prompts, chosen due to cost constraints; accordingly, the Gemma-based approach beats the rest in most cases for input length 8.
@@
The secondary results with SSIM and MSE of predicted observations are shown in Tables V and VI; as mentioned earlier, these metrics are not relevant enough to the dynamics and safety predictions of safety-critical systems.
@@ V-A Foundation Models and Applications
Large language models can also complete complex sequence tasks that are not limited to text but also extend — without specialized training — to domains such as robotics [16]. Except for applying to the decision-making field, language-guided abstraction [17] can transfer high-level task descriptions into task-relevant state abstractions by using a pre-trained language model. Abstract meaning representation also plays an important role in LLM and can improve the performance of LLM [18]. Since the human and robotic systems may have different understandings of natural languages, human and robot representations alignment is conducted to fix this gap [19].
As the embedded structure of LLM is a generative pre-trained transformer it can be used for prediction. Some researchers have tested LLM’s forecasting ability with humans and it shows the the ability of LLM is approaching the human level [20]. Some new foundation models like Chronos [21] are specialized to execute prediction tasks. Language can also be the input of robotic systems [22] to be trained for end-to-end language-controlled systems like RT2 [23] which can make basic semantic reasoning to finish corresponding tasks. The trustworthiness of robotic techniques based on foundation models is still an area that has not been widely expanded [24]. Some novel simulator also adopts LLM to generate various scenes for traffic simulations [25].
Method Input length SSIM MSE
@@
TABLE V: Comparison of Structural similarity index measure (SSIM) and mean square error (MSE) for cart pole’s predicted observations. Gray rows indicate the use of additional data..
Method Input length SSIM MSE
k=10 k=20 k=30 k=40 k=50 k=60 k=10 k=20 k=30 k=40 k=50 k=60
TABLE VI: Comparison of Structural similarity index measure (SSIM) and mean square error (MSE) for lunar lander’s predicted observations. Gray rows indicate the use of additional data.
V-B Safety Assurance of Autonomous Systems
The forecasting of trajectories occupies a pivotal role within autonomous systems, serving as a cornerstone of safety assurance. For sensors operating in high-dimensional spaces, integrating physics models is crucial for enhancing predictive accuracy. Works such as the Social ODE  [26] and Deep Kinematic Models  [27] exemplify this approach by merging deep learning techniques with physical models to address issues related to data unreliability. Another safety assurance method is to conduct the reachability analysis. Given the challenge of finding all reachable states in the non-linear dynamics and neural network controllers, verification tools [28, 29] are designed to develop accurate overapproximations of these reachable sets. However, these tools cannot work on high-dimensional controllers such as image-input controllers. High-dimensional verification processes could be tailored, for instance, through the application of generative models [30], and by transforming high-dimensional systems into their lower-dimensional counterparts [31]. Conformal prediction is also widely recognized for establishing boundaries around time series forecasts within a defined confidence level  [32, 33], which could be used to provide safety bounds for the trajectory with a specific confidence.
@@ VI Discussion
While the foundational world model demonstrates improved performance in certain aspects and produces more meaningful states, its dynamic predictions essentially rely on statistical correlations. Also, deploying foundation world models may require human judgment about the relevance of segmented classes; however, for many robotic systems, this effort is dwarfed by the significant data collection required to deploy standard world models or other supervised methods.

### v2

@@ IV-A Experimental Setup
TABLE IV: Safety prediction performance for the cart pole: F1 score and false positive rate. Gray rows indicate the use of additional data.
@@ IV-B Experimental Results
First, we discuss the state prediction results. For the cart pole, in order to show the safety-related state more intuitively, we calculate the angle by the centroids of the cart and pole to substitute the CD of the pole. As depicted in Table I and II, the errors of all methods are relatively low on short prediction horizons. As the value of $k$ increases, the error generally increases, indicating that predicting states further into the future is more difficult. Specifically, GPT 3.5 and Gemma alternately exhibit relatively low errors in most settings, especially when predicting the falling status. Table I shows the error for the lunar lander, where the foundation models do not perform as well as supervised learning methods on short inputs. Our investigation suggests that this shortcoming is due to insufficiently long input prompts, chosen due to cost constraints; accordingly, the Gemma-based approach beats the rest in most cases for input length 8.
@@
The secondary results with SSIM and MSE of predicted observations are shown in Tables V and VI. As mentioned earlier, these metrics are not relevant enough to the dynamics and safety predictions of safety-critical systems. Combining with the metric of the state prediction, it shows that he standard world model can only do well in the MSE and fails in SSIM, which means it has a low quality of predicting and reconstructing the observations.
Method Input length SSIM $\uparrow$ MSE $\downarrow$
k=10 k=20 k=30 k=40 k=50 k=60 k=10 k=20 k=30 k=40 k=50 k=60
TABLE V: Comparison of Structural similarity index measure (SSIM) and mean square error (MSE) for lunar lander’s predicted observations. Gray rows indicate the use of additional data.
@@ V-A Foundation Models and Applications
Large language models can also complete complex sequence tasks that are not limited to text but extend — without specialized training — to domains such as robotics [17]. In addition to applying to decision-making, language-guided abstraction [18] can transfer high-level task descriptions into task-relevant state abstractions by using a pre-trained language model. Abstract meaning representation also plays an important role in LLMs and can improve their performance [19]. Since human and robotic systems may have different understandings of natural languages, human and robot representation alignment is performed to close this gap [20].
As the embedded structure of LLM is a generative pre-trained transformer, it can be used for prediction. Some researchers have tested LLM’s forecasting ability with humans and it shows that the ability of LLM is approaching the human level [21]. Some new foundation models like Chronos [22] are specialized to execute non-dynamical prediction tasks like traffic, weather, and energy consumption. Language can also be the input of robotic systems [23] to be trained for end-to-end language-controlled systems like RT2 [24], which can perform basic semantic reasoning to finish its tasks. The trustworthiness of robotic techniques based on foundation models is still an area that has not been widely explored [25]. Besides applying the foundation model to control and forecast, some novel simulator also adopts LLM to generate various scenes for traffic simulations [26].
V-B Safety Assurance of Autonomous Systems
Trajectory forecasting plays a key role within autonomous systems for a variety of purposes, including safety assurance. For sensors operating in high-dimensional spaces, integrating physics models is crucial for enhancing predictive accuracy. Works such as the Social ODE [27] and Deep Kinematic Models [28] exemplify this approach by merging deep learning techniques with physical models to address issues related to data unreliability. Another popular safety assurance method is reachability analysis. Given the challenge of finding all reachable states in non-linear dynamics and neural network controllers, specialized verification tools [29, 30] are designed to develop accurate overapproximations of these reachable sets. However, these tools cannot work on high-dimensional controllers such as image-input controllers. High-dimensional verification processes could be tailored, for instance, through the application of generative models [31], and by approximating high-dimensional systems with their lower-dimensional counterparts [32]. Finally, conformal prediction is also widely recognized for establishing boundaries around time series forecasts within a defined confidence level  [33, 34], which could be used to provide safety bounds for the trajectory with a specific confidence.
Method Input length SSIM $\uparrow$ MSE $\downarrow$
@@ V-B Safety Assurance of Autonomous Systems
TABLE VI: Comparison of Structural similarity index measure (SSIM) and mean square error (MSE) for cart pole’s predicted observations. Gray rows indicate the use of additional data..
@@ VI Discussion
While our foundation world model demonstrates improved performance in certain aspects and produces physically meaningful states, its dynamic predictions essentially rely on statistical correlations. Also, deploying foundation world models may require human judgment about the relevance of segmented classes; however, for many robotic systems, this effort is dwarfed by the significant data collection required to deploy standard world models or other supervised methods.